Given the following information, what is the variance of a portfolio that is invested 25 percent in both stocks A and C, and 50 percent in stock B?

State of Economy Probability of State of Economy Rate of Return if State Occurs
Stock A Stock B Stock C
Boom 0.45 0.19 0.09 0.06
Natural 0.55 0.11 0.05 0.26
Answer #1

Return in boom =25%*0.19+50%*0.09+25%*0.06 =10.75%
Return in Natural =25%*0.11+50%*0.05+25%*0.26 =11.75%
Expected Return of Portfolio =0.45*10.75%+0.55*11.75% =11.30%

Variance of Portfolio =0.45*(10.75%-11.30%)^2+0.55*(11.75%-11.30%)^2 =0.002475% or 0.0025%